# Scaling Laws for Task-Stratified Knowledge in Post-Training Quantized Large Language Models

### v1

@@ Abstract
Large language models (LLMs) present significant deployment challenges due to their scale, with post-training quantization (PTQ) emerging as a practical compression solution. However, a comprehensive understanding of how PTQ precisely impacts diverse LLM knowledge capabilities remains elusive, and existing scaling laws for quantized models often overlook crucial PTQ-specific parameters and task-specific sensitivities. This paper addresses these gaps by conducting an extensive empirical investigation to establish task-stratified scaling laws. We disentangle LLM knowledge into memorization and utilization capabilities and develop a unified quantitative framework that incorporates model size, effective bit-width, calibration set size, and group size. Our central finding reveals that knowledge memorization exhibits markedly greater sensitivity to variations in effective bit-width, calibration set size, and model size compared to the more robust knowledge utilization. These findings offer a fine-grained understanding of PTQ’s impact and provide guidance for developing knowledge-aware quantization strategies that can better preserve targeted cognitive functions.
@@
Chenxi Zhou1,2, Pengfei Cao2,3, Jiang Li4, Jun Zhao2,3, Kang Liu2,3thanks: Corresponding Author 1School of Advanced Interdisciplinary Sciences, University of Chinese Academy of Sciences 2The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3School of Artificial Intelligence, University of Chinese Academy of Sciences 4College of Computer Science, Inner Mongolia University
@@ 3.3.1 Functional Form of the Task-Stratified Scaling Law
where $C_{task}$ is a task-specific constant scaling coefficient. $N$, $G$, $C_{b}$, and $B_{eff}$ represent the model size, calibration set size, group size, and effective bit-width, respectively, as previously defined. The exponents $\gamma_{task},\,\alpha_{task},\,\beta_{task},\,\delta_{task}$ are task-specific scaling parameters, quantifying the sensitivity of performance on that task type to each respective factor.

### v2

@@ Abstract
Large language models (LLMs) present significant deployment challenges due to their scale, with post-training quantization (PTQ) emerging as a practical compression solution. However, a comprehensive understanding of how PTQ precisely impacts diverse LLM knowledge capabilities remains elusive, and existing scaling laws for quantized models often overlook crucial PTQ-specific parameters and task-specific sensitivities. This paper addresses these gaps by conducting an extensive empirical investigation to establish task-stratified scaling laws. We disentangle LLM knowledge into memorization and utilization capabilities and develop a unified quantitative framework that incorporates model size, effective bit-width, calibration set size, and group size. Our central finding reveals that knowledge memorization exhibits markaredly greater sensitivity to variations in effective bit-width, calibration set size, and model size compared to the more robust knowledge utilization. These findings offer a fine-grained understanding of PTQ’s impact and provide guidance for developing knowledge-aware quantization strategies that can better preserve targeted cognitive functions.
@@
Chenxi Zhou1,2, Pengfei Cao2,3, Jiang Li4, Jun Zhao2,3, Kang Liu2,3thanks: Corresponding Author 1School of Advanced Interdisciplinary Sciences, University of Chinese Academy of Sciences 2The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3School of Artificial Intelligence, University of Chinese Academy of Sciences 4College of Computer Science, Inner Mongolia University zhouchenxi2025@ia.ac.cn, {pengfei.cao, jzhao, kliu}@nlpr.ia.ac.cn
@@ 3.3.1 Functional Form of the Task-Stratified Scaling Law
where $C_{task}$ is a task-specific constant scaling coefficient. $N$, $C_{b}$, $G$, and $B_{eff}$ represent the model size, calibration set size, group size, and effective bit-width, respectively, as previously defined. The exponents $\alpha_{task},\,\beta_{task},\,\gamma_{task},\,\delta_{task}$ are task-specific scaling parameters, quantifying the sensitivity of performance on that task type to each respective factor.